224. The scores of 40,000 county employees who took a recent civil service paper were normally distributed with μ = 70 and σ = 10.

The percentile rank of the score 85 is:
Answer #1

Step 1: Sketch the curve. The probability that X < 85 is equal to the blue area under the curve. 0.050 0.045 0.040 0.035 0.03

Step 2: Since u = 70 and o= 10 we have: X - u P(X < 85 ) = P(X – u < 85 – 70 ) = P (4. 85 – 70 10 Since = and 85770 = = 1.5 w